A mathematics textbook editor spent 7.5 hours making telephone calls, writing e-mails, and attending meetings. She spent twice as much time attending meetings as making telephone calls and 0.5 hours longer writing e-mails than making telephone calls. How many hours did she spend on each task?
Answer by stanbon(75887)

A mathematics textbook editor spent 7.5 hours making telephone calls, writing e-mails, and attending meetings. She spent twice as much time attending meetings as making telephone calls and 0.5 hours longer writing e-mails than making telephone calls. How many hours did she spend on each task?
----------------------
Equations:
t + w + a = 7.5 hrs.
a = 2t
w = t + 0.5
----------------------
Substitute and solve for "t"::
t + t+0.5 + 2t = 7.5
4t + 0.5 = 7.5
4t = 7
t = 7/4 hr = 1 hr 45 min (time for tele calls)
a = 2t = 3.5 hr (time attending meetings)
w = t+0.5 = 2hr 15 min
----------------------------
Cheers,
Stan H.
----------------